OpenStudy (lmvillamagua):

Hello guys! can you help me to solve the following problem. Given that f(x)=Sin(x) prove that f(x+2h)-f(x)=2Cos(x+h)Sin(h)

OpenStudy (er.mohd.amir):

since f(x)=sin x so f(x+2)=sin(x+2h) now f(x+2)- f(x)=sin (x+2h)-sinx use sin c-sin d=2 sin (c+d)/2 sin (c-d)/2

OpenStudy (lmvillamagua):

Ahhh ok ok. That makes a lot of sense. Thanks mate. A medal for you :)